### v1

# Distributed Threat Intelligence at the Edge Devices: A Large Language Model-Driven Approach

Abstract
With the proliferation of edge devices, there is a significant increase in attack surface on these devices. The decentralized deployment of threat intelligence on edge devices, coupled with adaptive machine learning techniques such as the in-context learning feature of large language models (LLMs), represents a promising paradigm for enhancing cybersecurity on low-powered edge devices. This approach involves the deployment of lightweight machine learning models directly onto edge devices to analyze local data streams, such as network traffic and system logs, in real-time. Additionally, distributing computational tasks to an edge server reduces latency and improves responsiveness while also enhancing privacy by processing sensitive data locally. LLM servers can enable these edge servers to autonomously adapt to evolving threats and attack patterns, continuously updating their models to improve detection accuracy and reduce false positives. Furthermore, collaborative learning mechanisms facilitate peer-to-peer secure and trustworthy knowledge sharing among edge devices, enhancing the collective intelligence of the network and enabling dynamic threat mitigation measures such as device quarantine in response to detected anomalies. The scalability and flexibility of this approach make it well-suited for diverse and evolving network environments, as edge devices only send suspicious information such as network traffic and system log changes, offering a resilient and efficient solution to combat emerging cyber threats at the network edge. Thus, our proposed framework can improve edge computing security by providing better security in cyber threat detection and mitigation by isolating the edge devices from the network.
Index Terms:
Edge Computing, Threat Intelligence, Machine Learning (ML), Large Language Model (LLM).I Introduction
Edge computing is ubiquitous from a privacy and security perspective [1]. We propose this approach with the aim of enhancing security by identifying and mitigating potential threats with the help of the Large Language Model (LLM) server. Large Language Model (LLM) refers to a subset of AI models that are optimized for specific tasks, such as natural language understanding and generation, which can further augment domain-specific contexts like personalized assistants [2] and threat detection [3], etc. The feature of in-context learning in the context of LLM can leverage the natural language understanding of LLM’s capabilities to learn new language skills within relevant and authentic linguistic contexts. OpenAI Generative Pre-trained Transformer (GPT) is a specific implementation of a LLM based on the transformer architecture, which has revolutionized this approach with billions of parameters. For instance, GPT-3.5 has 175 billion parameters compared to the previous version of GPT. Threat intelligence at the edge refers to the practice of gathering, analyzing, and applying threat intelligence data at the periphery of a network or system, where interactions with external entities occur. We propose a noble approach to edge threat intelligence at the edge that involves deploying lightweight AI models directly onto edge devices, such as routers, firewalls, and IoT devices. These models can continuously observe network traffic, system logs and configurations to identify patterns indicative of potential security threats.
II System Architecture
The key components of this approach to threat intelligence at the edge include a lightweight ML model trained to detect anomalies by analyzing network traffic and system logs, identify malicious activities, classify security threats in real-time, and communicate with the edge server by Message Queuing Telemetry Transport (MQTT) protocol, a lightweight publish-subscribe protocol to transfer data to the edge server and also among other edge devices and local edge, and a central LLM server trained with updated threat repositories. The central server can also update its knowledge through the in-context feature of GPT. The local edge server enables edge devices, and the Ml model provides edge devices with limited processing power and memory without compromising performance or resource efficiency. In figure 1, we describe the entire process into four main components:
1.
Edge devices with lightweight ML model: These ML models are designed to operate efficiently on edge devices with limited computational resources, enabling real-time threat analysis and response without relying heavily on cloud-based resources. By processing threat intelligence data locally on edge devices, it can provide real-time detection of security threats analyzing network data, device logs and other parameters without introducing significant delays or latency. This allows for immediate response actions, such as blocking malicious traffic or alerting the edge server, to mitigate potential risks. This model can vary from device to device, as every device has its own thread landscape in its deployment environment.
2.
MQTT which integrates edge devices with edge servers: Edge devices will be connected to the edge server via the MQTT channel. Also, the edge devices can communicate with each other to share data with the help of this channel, thus creating one-to-many communication.
3.
Edge server deployed locally: The edge server provides the MQTT queue, where edge devices basically exchange data with each other and the local edge server, preserving user privacy and security without transmitting sensitive information to external cloud servers. The edge server monitors the activity and, more specifically, the warning given by the compromised edge device by alerting the system administrator and blocking the edge device from communicating with others. This decentralized approach minimizes the risk of unauthorized modifications or data breaches in two way verification. Additionally, the communication latency is minimal as, in most scenarios, the edge server and edge devices are geographically located close together, which is critical for such communication.
4.
LLM server: LLM-centric central threat intelligence solution can adapt to and learn from new threats and attack vectors over time. Through continuous training and updates, LLM can improve its accuracy and effectiveness in identifying emerging security threats at the edge. In cases of unknown threats, the edge server, connected through a high-speed network, can communicate with the central LLM server about the type of attack, possible vulnerability, and effective solution for mitigating the risk in the shortest possible time.
This approach to threat intelligence at the edge can offer a proactive and efficient means of enhancing security in distributed computing environments, where traditional centralized threat detection mechanisms may be less effective. By leveraging lightweight AI models directly on edge devices, we can detect compromised edge devices with evolving cyber threats.
III Methodology
To protect the other edge devices from compromised devices in near real-time data involves multiple steps. Firstly, we assume the edge devices have network connectivity, and those are connected to the edge server via the MQTT protocol. All communication are encrypted with Secure Socket Layer (SSL) to prevent additional man-in-the-middle attacks. After receiving any messages from edge devices via the MQTT queue, the edge server analyzes the issue using a trained ML model. In an unknown case, edge server informs the central LLM server at a certain interval. In order to identify any suspicious activity, the lightweight AI model implemented with TensorFlow Lite [4] can detect. After detection with other meta-information like location, severity, etc., it will alert the edge server. If the edge server finds that request, analyze it and respond to other edge devices, notifying the monitoring team. In such an environment, other edge devices will not continue to communicate with the infected devices protecting the valuable data from the attacker.
For experimentation, we chose two raspberry Pis and one android mobile phone. For edge intelligence, we plan to deploy a trained tensorflow model to the rasberry Pi and convert the tensorflow ML model to the tensorflow Lite deployed on a cell phone or any wide range of IoT chips and microcontrollers. As every device at the edge gets a ML model, it becomes an intelligent device. Also, we use an edge server with an Intel (R) CoreTM i7 and 64 GB of RAM, which has an MQTT server running. All the edge devices are connected to the MQTT edge server via a client application. Furthermore, the edge server is connected to a central LLM server, which we consider OpenAI GPT 3.5 Turbo API which we will train with popular threat libraries. Moreover, We plan to enhance central server intelligence through in-context learning of GPT where edge servers can update the server model through the central LLM via a REST API call.
IV Conclusion and Future Work
The innovative side of this approach lies in the integration of LLM and diverse other popular technologies like ML application, edge server intelligence, MQTT to tackle cybersecurity challenges specific to lightweight edge devices. The proposed framework will have a relative impact on the field of edge device security by addressing wider white box attacks like zero days attack, poisoning attack etc. considering resource limitations while maintaining data privacy and scalability. In the future, we will provide a practical demonstration of the system to demonstrate the framework’s applicability in real-world deployment.